Through: (-2,3), slope= 1/2
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

point slope y-y1=m(x-x1), m slope (x1, y1) point
y-3=(1/2)(x+2)
y=(1/2)x+4

Answer by ikleyn(52908)